# Mutual Information Free Topological Generalization Bounds via Stability

### v1

@@ Abstract
Providing generalization guarantees for stochastic optimization algorithms is a major challenge in modern learning theory. Recently, several studies highlighted the impact of the geometry of training trajectories on the generalization error, both theoretically and empirically. Among these works, a series of topological generalization bounds have been proposed, relating the generalization error to notions of topological complexity that stem from topological data analysis (TDA). Despite their empirical success, these bounds rely on intricate information-theoretic (IT) terms that can be bounded in specific cases but remain intractable for practical algorithms (such as ADAM), potentially reducing the relevance of the derived bounds. In this paper, we seek to formulate comprehensive and interpretable topological generalization bounds free of intractable mutual information terms. To this end, we introduce a novel learning theoretic framework that departs from the existing strategies via proof techniques rooted in algorithmic stability. By extending an existing notion of hypothesis set stability, to trajectory stability, we prove that the generalization error of trajectory-stable algorithms can be upper bounded in terms of (i) TDA quantities describing the complexity of the trajectory of the optimizer in the parameter space, and (ii) the trajectory stability parameter of the algorithm. Through a series of experimental evaluations, we demonstrate that the TDA terms in the bound are of great importance, especially as the number of training samples grows. This ultimately forms an explanation of the empirical success of the topological generalization bounds.
@@ 1 Introduction
The generalization capabilities of deep learning models play a crucial role in our understanding of modern machine learning. Explaining generalization and, thus, the success of current neural networks is still an open problem and an active research area [ZBH+17, ZBH+21]. Following classical statistical learning theory [SSBD14], supervised learning tasks can be framed as a population risk minimization problem, defined as:
$\min_{w\in\mathbb{R}^{d}}\left\{\mathcal{R}(w):=\mathbb{E}_{z\sim\mu_{z}}[\ell(w,z)]\right\},$ (1)
@@
where $z\in\mathcal{Z}$ denotes the data, distributed according to a probability distribution $\mu_{z}$ on the data space $\mathcal{Z}$ and $\ell:\mathbb{R}^{d}\times\mathcal{Z}\to\mathbb{R}$, the composed loss function. In practice, we have $\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$, where $\mathcal{X}$ denotes a feature space and $\mathcal{Y}$ is a label space. Since the data distribution $\mu_{z}$ is unknown, it is in general not possible to directly solve (1); instead, we typically solve the empirical risk minimization problem (2): for $n$ given data points $S:=(z_{1},\dots,z_{n})\sim\mu_{z}^{\otimes n}:=\mu_{z}\otimes\dots\otimes\mu_{z}$ independently sampled from the data distribution $\mu$, we consider:
$\min_{w\in\mathbb{R}^{d}}\left\{\widehat{\mathcal{R}}_{S}(w):=\frac{1}{n}\sum_{i=1}^{n}\ell(w,z_{i})\right\}$ (2)
In most applications, the problem (2) is solved by the means of stochastic optimization algorithms, such as stochastic gradient descent (SGD) or ADAM [KB14]. These algorithms often take an iterative form and we denote $w_{k}$ the parameter that is learned after $k$ iterations.
@@
Recent studies suggest that the training trajectory, i.e., the set of iterates $w_{k}$ obtained by the algorithm, may be paramount in explaining the generalization ability of modern machine learning models [XTYY24, FZY+24, ADS+24]. Indeed, since the empirical risk (2) can contain many (typically infinitely many) local minima with different generalization properties, the trajectories of the optimization algorithm near a local minimum offer a concise way to assess the quality of the local minimum: they inherently reflect the influence of the optimization algorithm, the choice of hyperparameters, and the dataset—factors that are all essential for deriving meaningful generalization bounds.
More precisely, the trajectory of the optimizer is essentially a random subset of $\mathbb{R}^{d}$, that is defined as follows: $\mathcal{W}_{S,U}:=\{w_{t}:t_{0}\leq t\leq T\}\subset\mathbb{R}^{d}$, where $t_{0}$ is chosen such that $w_{t_{0}}$ is near a local minimum that is found by the optimizer, say $w_{\star}$, $U$ represents the algorithmic randomness (e.g., batch indices), which is independent of the other random variables. In essence, $\mathcal{W}_{S,U}$ is the part of the basin corresponding to $w_{\star}$ that is accessible to the optimization algorithm. To measure the quality of the local minimum $w_{\star}$ by using the structure of the trajectory $\mathcal{W}_{S,U}$, the main quantity of interest has been the worst-case generalization error over $\mathcal{W}_{S,U}$ , defined as follows:
$G_{S}(\mathcal{W}_{S,U}):=\sup_{w\in\mathcal{W}_{S,U}}\left(\mathcal{R}(w)-\hat{\mathcal{R}}_{S}(w)\right).$ (3)
Upper-bounding $G_{S}(\mathcal{W}_{S,U})$ requires uniform convergence results on the random, data-dependent hypothesis set $\mathcal{W}_{S,U}$, which is a notoriously challenging problem [FGK+19, SvEH+23, DVDS24].
In this context, recent studies [SSDE20, DDS23, ALRS23, HSKM22, DVDS24, BLGS21, ADS+24] have established upper-bounds on (3) and drawn rigorous links between the topological structure of $\mathcal{W}_{S,U}$ and $G_{S}(\mathcal{W}_{S,U})$. These bounds are of the following form [^1] We use the notation $\lesssim$ in informal statements where absolute constants have been omitted.:
@@
$G_{S}(\mathcal{W}_{S,U})\lesssim\sqrt{\frac{\mathbf{C}(\mathcal{W}_{S,U})+\mathrm{IT}+\log(1/\zeta)}{n}},$ (4)
with probability at least $1-\zeta$. The bound in (4) has two main ingredients: (i) The first term $\mathbf{C}(\mathcal{W}_{S})$, which we call the topological complexity in accordance with [ADS+24], is a geometric quantity of the trajectory, measuring some notion of complexity of $\mathcal{W}_{S}$. Several topological complexities have been proposed, inspired by fractal geometry [SSDE20, DDS23, DV23] and topological data analysis (TDA) [BLGS21, ALRS23, ADS+24]. (ii) The second term, $\mathrm{IT}$, is an information-theoretic (IT) term, typically the total mutual information between the dataset $S$ and the trajectory $\mathcal{W}_{S}$; measuring the statistical dependence between $S$ and $\mathcal{W}_{S,U}$ [DVDS24].
The topological complexity term $\mathbf{C}(\mathcal{W}_{S,U})$ appearing in these bounds are both computable and practically informative. In contrast, the $\mathrm{IT}$ term remains opaque and computationally intractable; its behavior with respect to the sample size $n$ is not even well understood in the general case [DVDS24]. As a result, prior studies have adopted a heuristic approach by entirely discarding the IT term in their empirical evaluations and concentrating solely on $\mathbf{C}(\mathcal{W}_{S,U})$. These studies reported striking empirical results, demonstrating a strong correlation between the generalization error and the proposed topological complexity measures (an observation we will also report in our experiments) [^2] While [TGRW+24] recently raised some concerns about the practical validity of certain fractal dimension–based bounds, [ADS+24] showed that specific topological descriptors—such as positive magnitude and $\alpha$-weighted lifetime sums, which we will detail in the next section, consistently yield strong empirical performance across a wide range of modern deep learning scenarios.. However, the exclusion of the IT term renders such evaluations theoretically incomplete, and the underlying reason for the empirical success of the topological complexity measures remains unclear.
To address this issue, an initial step was taken in [SvEH+23], where the authors introduced the notion of ‘algorithm-dependent Rademacher complexity’ (ARC) and derived generalization bounds with a topological flavor, notably avoiding the $\mathrm{IT}$ term. However, ARC is based on combining parameters in a combinatorial fashion (similar to the approach that has been used for conditional mutual information-based bounds [SZ20]). As a result, the associated topological quantities were not tied to the data-dependent optimization trajectory $\mathcal{W}_{S,U}$, but rather to an abstract, theoretically constructed hypothesis set. Consequently, while the bounds are insightful from a theoretical standpoint, they fall short of explaining why $\mathbf{C}(\mathcal{W}_{S,U})$ proves effective in practice.
In this study, our goal is to bridge the gap between topological generalization bounds and the aforementioned empirical observations, by deriving generalization bounds that are directly expressed in terms of the $\mathbf{C}(\mathcal{W}_{S,U})$ while entirely avoiding the information-theoretic terms.
@@
Our main contributions are summarized below.
  • Inspired by [FGK+19], we define the notion of trajectory stability, that is adapted to stochastic learning algorithms, i.e., producing random trajectories. Based on this stability condition, we prove expected topological generalization bounds of the form:
@@
    $\mathbb{E}\left[\sup_{w\in\mathcal{W}_{S,U}}\left(\mathcal{R}(w)-\widehat{\mathcal{R}}_{S}(w)\right)\right]\lesssim\beta_{n}^{1/3}\left(1+\mathbb{E}\left[\sqrt{\mathbf{C}(\mathcal{W}_{S,U})}\right]\right),$
    where $\beta_{n}$ is the stability parameter and $\mathbf{C}(\mathcal{W}_{S,U})$ is a topological complexity identical to some of the TDA-inspired terms used in [BLGS21, ADS+24]. In particular, we recover the so-called $\alpha$-weighted lifetime sums and positive magnitude based on the Euclidean distance (see Section 2).
  • We further investigate our new trajectory stability condition and show that a variant of the classical notion of algorithmic stability [HRS16] implies trajectory stability. Hence, if the optimization algorithm is algorithmically stable in a more classical sense, existing algorithmic stability results imply that, in various cases, the trajectory stability parameter $\beta_{n}$ will be of order $\mathcal{O}(1/n)$, rendering our bounds $\mathcal{O}(n^{-1/3}\sqrt{\mathbf{C}(\mathcal{W}_{S})})$.
  • We extend recent empirical studies on topological generalization bounds and explain the interplay between stability and topology, emphasizing the role of sample size as a key factor in the analysis of topological generalization bounds.
Our results demonstrate that, if the optimizer is trajectory-stable, the information-theoretic terms in the bound Equation 4 can be safely omitted, at the expense of a reduced convergence rate [^3] A similar convergence rate was also reported in [DDS23] in an attempt to reduce the impact of the IT terms., i.e., from $n^{-1/2}$ to $n^{-1/3}$. We believe that our findings reduce the gap between theory and practice, paving the way for even more practical topological generalization bounds with new empirical insights for large-scale datasets. The implementation will be made publicly available at https://github.com/bastianlb/MI-free-topo-bounds. All the proofs are provided in the appendix.
2 Preliminaries Technical Background
We now introduce technical background regarding the topological complexities used in this work (Section 2.1) and algorithmic stability (Section 2.2).
2.1 Topological Complexity Measures
$\alpha$-Weighted lifetime sum
The $\alpha$-weighted lifetime sum is one of the building blocks of the persistent homology dimension [Sch20, BLGS21]. Lifetime intervals encode how long topological features (like connected components) persist across scales [BGND+19, HKNU17, HKND19]. Informally, it tracks a finite set’s “connected components” at different scales. For simplicity and readability, we provide an equivalent formulation based on minimum spanning trees [Sch20, KLS06].
Let $X$ be a finite set of points endowed with a metric $\rho$. A tree over X is a connected acyclic undirected graph whose vertices are the points in X. Given an edge $e$ linking the points a and b, we define its cost as $|e|:=\rho(a,b)$ The $\alpha$-weighted lifetime sums are defined as:
$\forall\alpha\geq 0,\quad\textbf{E}_{\rho}^{\alpha}(X):=\sum_{e\in\mathrm{MST}(X)}|e|^{\alpha},$
where $\mathrm{MST}(X)$ denotes a minimum spanning tree over the graph defined by $X$.
Positive magnitude
The concept of metric magnitude was introduced in [Lei13], it is critical in geometric measure theory and integral geometry [Mec13]. In [ADS+24] a variant called positive magnitude was introduced and rigorously linked to generalization error. Informally, (positive) magnitude can be interpreted as the effective number of distinct points in a space at a given scale. We now give the definition of positive magnitude. For a “scale” parameter $s>0$ and a finite metric space $(X,\rho)$, we define a weighting of the scaled space $(s\cdot X,\rho)$ as a vector $\gamma_{s}:X\to\mathbb{R}$ such that for any $a\in X$, we have $\sum_{b\in X}e^{-s\rho(a,b)}\gamma_{s}(b)=1$. The existence of such weighting is ensured in the finite Euclidean setting [Mec15]. The positive magnitude of $(s\cdot X,\rho)$ is then defined as:
$\mathbf{PMag}_{\rho}(s\cdot X):=\sum_{b\in X}\gamma_{s}(b)_{+},$ (5)
where $\gamma_{+}:=\max\{0,\gamma\}$. The parameter $s$ can be seen as a scale through which we look at $(X,\rho)$.
2.2 Algorithmic stability
Algorithmic stability is a fundamental concept in learning theory [BE02, BKZ20]. It has proven successful in providing generalization guarantees for widely used stochastic optimization algorithms [BFGT20, FV19, ZGRS23, HRS16]. In this paper, we consider a variant of algorithmic stability, called uniform argument stability, introduced by [BFGT20].
In all the following, we say that two datasets $S,S^{\prime}\in\mathcal{Z}^{n}$ are neighboring datasets if they differ by one element, which we denote $S\simeq S^{\prime}$.
Definition 2.1 (Uniform argument stability [BFGT20]).

### v2

# Stability, Complexity and Data-Dependent Worst-Case Generalization Bounds

@@ Abstract
Providing generalization guarantees for stochastic optimization algorithms remains a key challenge in learning theory. Recently, numerous works demonstrated the impact of the geometric properties of optimization trajectories on generalization performance. These works propose worst-case generalization bounds in terms of various notions of intrinsic dimension and/or topological complexity, which were found to empirically correlate with the generalization error. However, most of these approaches involve intractable mutual information terms, which limit a full understanding of the bounds. In contrast, some authors built on algorithmic stability to obtain worst-case bounds involving geometric quantities of a combinatorial nature, which are impractical to compute. In this paper, we address these limitations by combining empirically relevant complexity measures with a framework that avoids intractable quantities. To this end, we introduce the concept of random set stability, tailored for the data-dependent random sets produced by stochastic optimization algorithms. Within this framework, we show that the worst-case generalization error can be bounded in terms of (i) the random set stability parameter and (ii) empirically relevant, data- and algorithm-dependent complexity measures of the random set. Moreover, our framework improves existing topological generalization bounds by recovering previous complexity notions without relying on mutual information terms. Through a series of experiments in practically relevant settings, we validate our theory by evaluating the tightness of our bounds and the interplay between topological complexity and stability.
@@ 1 Introduction
Explaining the generalization capabilities of modern deep learning models is still an open problem and an active research area (Zhang et al., 2017, 2021). Classically, supervised learning tasks can be framed as a population risk minimization problem, defined as:
$\textstyle\min_{w\in\mathbb{R}^{d}}\left\{\mathcal{R}(w):=\mathbb{E}_{z\sim\mu_{z}}[\ell(w,z)]\right\},$ (1)
@@
where $z\in\mathcal{Z}$ denotes the data, following a probability distribution $\mu_{z}$ on the data space $\mathcal{Z}$, and $\ell:\mathbb{R}^{d}\times\mathcal{Z}\to\mathbb{R}$ is the loss function. In practice, we have $\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$, where $\mathcal{X}$ denotes a feature space and $\mathcal{Y}$ is a label space. Since $\mu_{z}$ is unknown, it is in general not possible to directly solve Equation˜1; instead, we typically solve the empirical risk minimization problem: for a dataset $S:=(z_{1},\dots,z_{n})\sim\mu_{z}^{\otimes n}$ sampled i.i.d. from $\mu_{z}$, it corresponds to
$\textstyle\min_{w\in\mathbb{R}^{d}}\big\{\widehat{\mathcal{R}}_{S}(w):=\frac{1}{n}\sum_{i=1}^{n}\ell(w,z_{i})\big\}.$ (2)
To evaluate the validity of this replacement, it is essential to quantify the discrepancy between the two objectives. Thus, a central quantity of interest in learning theory is the generalization error, defined as $G_{S}(w):=\mathcal{R}(w)-\widehat{\mathcal{R}}_{S}(w)$.
To quantify the generalization error, a large body of work analyzes the problem at the level of a single weight vector $w$ by modeling the learning algorithm as a mapping $(S,U)\mapsto w_{S,U}$, where $S\in\mathcal{Z}^{n}$ represents the dataset and $U$ captures the algorithmic randomness. This viewpoint underlies, for example, information-theoretic approaches (Xu and Raginsky, 2017; Steinke and Zakynthinou, 2020; Pensia et al., 2018), as well as the algorithmic stability frameworks (Bousquet and Elisseeff, 2002; Hardt et al., 2016; Feldman and Vondrak, 2019). While these approaches provide generalization guarantees in various settings, the choice of which iterate to analyze remains unclear in the absence of a stopping criterion, and consequently cannot capture the full behavior of the generalization error.
@@
In contrast to focusing on a single weight vector, recent work has highlighted the role of the parameter trajectory – the sequence of iterates generated by the learning algorithm for solving Equation˜2 – in explaining the generalization behavior of modern machine learning models (Xu et al., 2024; Fu et al., 2024; Andreeva et al., 2024). Indeed, since the empirical risk can contain many local minima, the trajectory surrounding a local minimum offers a concise way to assess its quality. The trajectory itself inherently reflects the influence of the optimization algorithm, the choice of hyperparameters, and the dataset, factors that are essential for deriving meaningful generalization bounds.
To capture the geometric information of the parameter trajectory, several authors have proposed analyzing the worst-case generalization error along the entire trajectory, rather than relying on a single weight vector. Due the dependence of the parameter trajectory on the random variables $S$ (dataset) and $U$ (algorithmic noise), numerous studies on worst-case generalization bounds adopted the formulation of data-dependent random sets (Simsekli et al., 2020; Dupuis et al., 2023; Andreeva et al., 2023). In line with these previous works, we formalize a learning algorithm as a mapping,
$\textstyle\mathcal{A}:\bigcup_{n\geq 1}\mathcal{Z}^{n}\times\Omega\longrightarrow\mathrm{CL}(\mathbb{R}^{d}):=\{\text{closed subsets in }\mathbb{R}^{d}\},\quad\mathcal{A}:(S,U)\longmapsto\mathcal{W}_{S,U},$ (3)
where $S\in\mathcal{Z}^{n}$ is the dataset and $U\in\Omega$ is an independent random variable representing the algorithm randomness. This formalization includes (but is not limited to) the following examples.
Example 1.1.
Consider a stochastic optimizer $w_{k+1}=F(w_{k},S,U)$ ($U$ is the batch noise) and $K_{1},K_{2}\in\mathds{N}^{\star}$. The parameter trajectory is defined as $\mathcal{A}(S,U)=\mathcal{W}_{S,U}:=\left\{w_{k},~K_{1}\leq k\leq K_{2}\right\}$.
Example 1.2.
Consider a dynamics of the form $\mathrm{d}W_{t}=-\nabla\widehat{\mathcal{R}}_{S}(W_{t})\mathrm{d}t+\sigma\mathrm{d}B_{t}$, where $U:=(B_{t})_{t\geq 0}$ is a Brownian motion. Let $T>0$, the parameter trajectory is $\mathcal{W}_{S,U}:=\left\{W_{t},~t\in[0,T]\right\}$.
The following Examples 1.3 and 1.4 illustrate that this formulation also includes classical learning-theoretic frameworks.
Example 1.3.
Singleton bounds correspond to the case where we choose $\mathcal{A}(S,U):=\{w_{S,U}\}$.
Example 1.4.
Classical worst-case (uniform) generalization error over a data-independent hypothesis set $\mathcal{W}$ (Shalev-Shwartz and Ben-David, 2014) correspond to a constant mapping $\mathcal{A}(S,U)=\mathcal{W}$.
In this general setting, the main quantity of interest is the worst-case generalization error over $\mathcal{W}_{S,U}$,
$\textstyle G_{S}(\mathcal{W}_{S,U}):=\sup_{w\in\mathcal{W}_{S,U}}\big(\mathcal{R}(w)-\widehat{\mathcal{R}}_{S}(w)\big).$ (4)
In the case of Example˜1.4, this problem has been classically analyzed through the Rademacher complexity over $\mathcal{W}$ (Bartlett and Mendelson, 2002). However, as noted by Dupuis et al. (2023), these classical techniques break down for data-dependent random sets.
To overcome this limitation, many studies have employed information-theoretic techniques, particularly within the “fractal-based” literature (Simsekli et al., 2020; Birdal et al., 2021). A unifying perspective recently emerged with the PAC-Bayesian theory for random sets (Dupuis et al., 2024), which was recently employed by Andreeva et al. (2024) to establish generalization bounds based on novel topological complexity measures. Informally, all these bounds are of the following form [^1] We use the notation $\lesssim$ in informal statements where absolute constants have been omitted.:
@@
${\textstyle\sup_{w\in\mathcal{W}_{S,U}}}\big(\mathcal{R}(w)-\widehat{\mathcal{R}}_{S}(w)\big)\lesssim\sqrt{\frac{\mathbf{C}(\mathcal{W}_{S,U})+\mathrm{IT}+\log(1/\zeta)}{n}},$ (5)
with probability at least $1-\zeta$. The term $\mathrm{IT}$ is an information-theoretic (IT) term, typically the total mutual information between the dataset $S$ and the set $\mathcal{W}_{S,U}$ The aforementioned bounds differ in the choice of complexity measure $\mathbf{C}(\mathcal{W}_{S,U})$, but all include an IT term. For concrete examples, we refer the reader to (Dupuis et al., 2023, 2024; Andreeva et al., 2024). The presence of such IT term is a major drawback of these studies, as it is computationally intractable and not well-understood in the general case (Dupuis et al., 2024), and they can potentially be infinite.
[…]
Figure 1: (Left) Evolution of the loss function $\ell(\cdot,Z)$ for a fixed sample $Z\in\mathcal{Z}$ over $T=1000$ iterations, for two neighboring datasets $S,S^{\prime}\in\mathcal{Z}^{n}$. While the classical notion of algorithmic stability measures the error at a specific iteration $t$, our stability notion extends this perspective to the entire training trajectory. (Right) Numerical estimation of our new random set stability parameter ($\beta_{n}$) as $n$ increases. The experiments demonstrate that $\beta_{n}$ decreases with larger $n$.
@@
In this study, we propose to address the aforementioned issues by introducing a novel framework for deriving worst-case generalization bounds on data-dependent random sets, which are empirically relevant and do not contain intractable mutual information terms. Inspired by Foster et al. (2019), we adopt a stability-based approach, which we adapt to exploit the algorithm randomness in an explicit way. Based on our new assumption, we show that $G_{S}(\mathcal{W}_{S,U})$ can be upper bounded in terms of the stability parameter and a Rademacher complexity term. This allows us to take advantage of the best of both worlds by (1) avoiding the intricate information-theoretic (IT) terms and (2) relating the generalization error to topologically meaningful complexity measures, hence, providing IT-terms free versions of the topological generalization bounds of Birdal et al. (2021); Andreeva et al. (2024). Our contributions are detailed below.
  • We introduce the notion of random set stability, specifically designed for stochastic learning algorithms. More precisely, we define a new stability parameter $\beta_{n}$ measuring how much the random set $\mathcal{W}_{S,U}$ changes if we replace one or more data points in $S$, in a sense that is formally introduced in ˜3.1 below.
    Moreover, we relate $\beta_{n}$ to classical stability notions, providing a systematic procedure to establish random set stability in diverse settings.
  • We derive an expected worst-case generalization bound in terms of a Rademacher complexity term evaluated over the empirically relevant set $\mathcal{W}_{S,U}$ and the stability parameter.
  • We apply our framework to obtain generalization bounds based on the fractal and topological complexity measures introduced in (Birdal et al., 2021; Andreeva et al., 2024) without any intractable mutual information terms. These bounds read as follows:
@@
    $\textstyle\mathbb{E}\left[\sup_{w\in\mathcal{W}_{S,U}}\left(\mathcal{R}(w)-\widehat{\mathcal{R}}_{S}(w)\right)\right]\lesssim\beta_{n}^{1/3}\left(1+\mathbb{E}\left[\sqrt{\log\mathbf{C}(\mathcal{W}_{S,U})}\right]\right),$
    where $\beta_{n}$ is the stability parameter and $\mathbf{C}(\mathcal{W}_{S,U})$ is the complexity measure (see Section˜4). Therefore, we provide the first fully computable topological bounds for practically used optimization algorithms, hence, addressing the main issues of the aforementioned previous works.
  • We provide a systematic empirical investigation of the tightness of our bounds.
  • We empirically investigate the previously introduced topological quantities and their interplay with stability as the sample size $n$ varies, extending the correlation analyses of prior studies, and therefore providing evidence supporting the validity of our stability-based approach.
Our implementation is publicly available at https://github.com/bastianlb/MI-free-topo-bounds. Detailed proofs for all theoretical results are provided in the Appendix.
2 Technical Background
In this section, we provide some technical background on algorithmic stability. One of our contributions, described in Section˜3.1, is to extend these notions to data-dependent random sets.
Algorithmic stability is a fundamental concept in learning theory (Bousquet and Elisseeff, 2002; Bousquet et al., 2020). It has proven successful in providing generalization guarantees for widely used stochastic optimization algorithms (Feldman and Vondrak, 2019; Zhu et al., 2023; Hardt et al., 2016). To position our work within existing stability-based bounds, we consider a variant of algorithmic stability, called uniform argument stability (Bassily et al., 2020). As with most stability notions, it is defined with respect to a single iteration. Therefore, the learning algorithm used in the next definition corresponds to the setting of Example˜1.3 and can be seen as a randomized mapping $(S,U)\longmapsto\mathcal{A}(S,U)\in\mathbb{R}^{d}$. In all the following, we say that two datasets $S,S^{\prime}\in\mathcal{Z}^{n}$ are neighboring datasets if they differ by one element, which we denote $S\simeq S^{\prime}$.
Definition 2.1.